On January 1, 2017, Shamrock Co. purchased 26,000 shares (a 10% interest) in Elton John Corp. for $1,380,000. At the time, the book value and the fair value of John’s net assets were $14,200,000.

On July 1, 2018, Shamrock paid $3,000,000 for 52,000 additional shares of John common stock, which represented a 20% investment in John. The fair value of John’s identifiable assets net of liabilities was equal to their carrying amount of $15,500,000. As a result of this transaction, Shamrock owns 30% of John and can exercise significant influence over John’s operating and financial policies. (Any excess fair value is attributed to goodwill.)

John reported the following net income and declared and paid the following dividends.

Net Income

Dividend per Share

Year ended 12/31/17 $750,000 None
Six months ended 6/30/18 480,000 None
Six months ended 12/31/18 807,000 $1.65


Determine the ending balance that Shamrock Co. should report as its investment in John Corp. at the end of 2018.

Investment in Elton John Corp. $
Answer #1

A computation of the ending balance that Shamrock Co. should report as its investment in John Corp. at the end of 2018. is as follows:

Investment in John Corp. 1/1/17 = $1,380,000
Additional purchase 7/1/18 = $3,000,000
Adjustment for 2017 income (prior period) ($750,000 * 10%) = $75,000
Adjustment for 2018 income to 6/30 (prior period) ($480,000 * 10%) = $48,000
Income (7/1/18–12/31/18) ($807,000 * 30%) = $242,100
Less: Dividends (7/1/18–12/31/18) (78,000 shares * $1.65) = $(128,700)
Investment in Elton John Corp. as on 12/31/18 = $4,616,400

Under GAAP, goodwill is not amortized.